# SpreadsheetLLM: Encoding Spreadsheets for Large Language Models

### v1

@@ Abstract
Spreadsheets are characterized by their extensive two-dimensional grids, flexible layouts, and varied formatting options, which pose significant challenges for large language models (LLMs). In response, we introduce SpreadsheetLLM, pioneering an efficient encoding method designed to unleash and optimize LLMs’ powerful understanding and reasoning capability on spreadsheets. Initially, we propose a vanilla serialization approach that incorporates cell addresses, values, and formats. However, this approach was limited by LLMs’ token constraints, making it impractical for most applications. To tackle this challenge, we develop SheetCompressor, an innovative encoding framework that compresses spreadsheets effectively for LLMs. It comprises three modules: structural-anchor-based compression, inverse index translation, and data-format-aware aggregation. It significantly improves performance in spreadsheet table detection task, outperforming the vanilla approach by 25.6% in GPT4’s in-context learning setting. Moreover, fine-tuned LLM with SheetCompressor has an average compression ratio of 25×, but achieves a state-of-the-art 78.9% F1 score, surpassing the best existing models by 12.3%. Finally, we propose Chain of Spreadsheet for downstream tasks of spreadsheet understanding and validate in a new and demanding spreadsheet QA task. We methodically leverage the inherent layout and structure of spreadsheets, demonstrating that SpreadsheetLLM is highly effective across a variety of spreadsheet tasks.
@@
Yuzhang Tianthanks:   Equal contribution.thanks:   Work during internship at Microsoft., Jianbo Zhao11footnotemark: 122footnotemark: 2, Haoyu Dongthanks:   Corresponding author., Junyu Xiong, Shiyu Xia, Mengyu Zhou, Yun Lin22footnotemark: 2, José Cambronero, Yeye He, Shi Han, Dongmei Zhang Microsoft Corporation
@@ 1 Introduction
We conducted a comprehensive evaluation of our method on a variety of LLMs. Our experiments show that SheetCompressor significantly reduces token usage for spreadsheet encoding by 96%. Moreover, SpreadsheetLLM has shown exceptional performance in spreadsheet table detection, the foundational task of spreadsheet understanding, surpassing the previous SOTA method by 12.3% Dong et al. (2019b). We also applied SpreadsheetLLM to a representative spreadsheet QA task. Inspired by the Chain of Thought (CoT) methodology Zheng et al. (2023); Jiang et al. (2023b), we propose Chain of Spreadsheet (CoS) to decompose spreadsheet reasoning into a table detection-match-reasoning pipeline. It significantly outperformed existing SOTA methods for table QA Herzig et al. (2020); Cheng et al. (2022). Our primary contributions are summarized as follows:
@@
  • We fine-tune a variety of cutting-edge LLMs to achieve optimal performance on spreadsheet table detection, and demonstrate the high effectiveness of SpreadsheetLLM in accurately understanding complex spreadsheet layouts and structures.
@@
Figure 2: Illustration of the SheetCompressor framework. The original spreadsheet contains two tables, featuring numerous data entries or hierarchical headers, which can be viewed in the supplementary materials. The completed spreadsheet consists of 576 rows and 23 columns, with an vanilla encoding of 61,240 tokens. Initially, we first extract cells using structural anchors, rearranging them into a smaller 24×8 sheet. Subsequently, we perform index-invert, removing empty cells. Finally, we aggregate cells based one data formats, achieving an extremely compact representation of the spreadsheet, which contains only 708 tokens.
@@ 2 Related Work
Spreadsheet representation involves converting the spreadsheets into specific representations for different models. There are various methods for spreadsheet (table) representation. Dong et al. (2019a, b) enhance Mask-RCNN to leverage spatial and visual information in spreadsheets, and Deng et al. (2024) explores the usage of LLMs to evaluate image tables, but it doesn’t work well for spreadsheet images as input to VLMs Xia et al. (2024). To capture sequential semantics in rows and columns, LSTMs are further adopted Nishida et al. (2017); Gol et al. (2019) in row&column directions. Pre-trained LMs Dong et al. (2022) are then proposed to understand spreadsheets Wang et al. (2021). Recent studies Zhang et al. (2023); Li et al. (2023b); Sui et al. (2023) have explored the efficacy of using Markdown and HTML for table representation. However, they are not well suited to spreadsheets due to their single table input, as experiments show in Appendix B.
@@
Spreadsheet understanding is enabling for a series of spreadsheet tasks, such as table question answering analysis He et al. (2024); Cheng et al. (2021b, 2022); Jiang et al. (2022); Herzig et al. (2020), table extraction Chen and Cafarella (2013, 2014); Li et al. (2024), formula or code generation Chen et al. (2021); Cheng et al. (2021a); Joshi et al. (2024); Chen et al. (2024); Li et al. (2023a), error detection Wang and He (2019); Dou et al. (2016), etc. In this paper, we choose spreadsheet QA, one of the most demanded spreadsheet analysis tasks. It is an extension of the Table QA task in spreadsheet data, with the additional complexity of detecting and matching multiple tables within a spreadsheet.
@@
Related work suggests that the performance of LLMs degrades significantly with long contexts Liu et al. (2024); Xu et al. (2023). Efforts to improve model performance and reduce costs have led to the development of compression techniques for long prompts. Some researchers employ information-theory metrics to filter out redundant information Li (2023); Jiang et al. (2023a). Additionally, specialized models have been proposed to optimize prompt compression Pan et al. (2024). However, these strategies primarily address natural language prompts and may not suit tabular data, potentially leading to considerable structure and data information loss. DBCopilot Wang et al. (2023) enables text-to-SQL conversion on large databases through schema routing. However, due to LLMs’ insufficient ability in understanding inherent multi-table layouts and complex table structures that cannot execute queries similar to SQL, schema routing is impractical, restricting the broader application of cutting-edge tabular works Cheng et al. (2022); Li et al. (2023b); Sui et al. (2024) on spreadsheet data.
@@ 3.2 Structural-anchor-based Extraction
where $r_{p}=\left\{Cell_{i,j}\right\}_{i=p,j\in n}$ and $c_{q}=\left\{Cell_{i,j}\right\}_{i\in m,j=q}$. Using these anchor points, our method discards rows and columns that are located more than $k$ units away from any anchor point, because they rarely serve as table boundaries. The parameter $k$ serves as a threshold to control the scope of neighborhood retention, effectively eliminating areas predominantly filled with homogeneous data that do not contribute to an understanding of the spreadsheet’s layout and structure. We explored the effects of different $k$ values in an ablation study, as detailed in Appendix D.1.
@@ 3.3 Inverted-index Translation
Spreadsheets often contain numerous empty rows, columns, and scattered cells. The standard encoding method, as detailed in Section 3.1, employs a grid-based method that pairs cell addresses with their contents. This approach necessitates recording empty cells to maintain the spreadsheet’s two-dimensional structure, which significantly increases token consumption. Furthermore, cells with identical values are encoded repeatedly, further exacerbating token usage.
@@ 3.4 Data-format-aware Aggregation
In spreadsheets, adjacent cells typically share the same data format. As shown in Figure 2 (3), column C records the sell-in billed revenue for different products. Nonetheless, the concrete numerical values are not essential for understanding the structure and semantics of the spreadsheet (although there might loss of fine-trained details of exact quantities, e.g., "18,476" and "18,674", this does not impact our comprehension that this column represents revenue). In contrast, the data type is critical for understanding spreadsheets. On one hand, data types represent fundamental semantic properties, such as "time" or "phone number". It motivates us to implement rules to match the value of the cell to different data types. On the other hand, in contrast to detailed numerical values, identical data types may be compressed through clustering, thereby reducing the number of tokens.
@@ 3.5 Chain of Spreadsheet
The query and the identified table section are re-input into the LLM. The model then processes this information to generate an accurate response to the query.
@@ 4.1.1 Dataset
We used the dataset introduced by Dong et al. (2019b), a benchmark dataset of real-world spreadsheets with annotated table boundaries. Due to the complexity and ambiguity of precise address labeling (the Fleiss Kappa on the test set is 0.830), we further implemented the quality improvement pipeline on the test set by five human professions, as detailed n in Appendix E. To this end, we obtained a highly validated test set containing 188 spreadsheets. Based on the token usage of the vanilla encoding method, we divided the test set into four categories: Small, Medium, Large, and Huge, with a partition of 64:32:70:22. More details are shown in Appendix F. We adopted the Error-of-Boundary 0 (EoB-0) metric for evaluation on 188 spreadsheets with 311 tables. EoB-0 requires exact match of the top, left, bottom, and right boundaries.
@@ 4.1.2 Experiment Setup
To evaluate the performance of SpreadsheetLLM, we chose TableSense-CNN Dong et al. (2019b) as the baseline due to its previously demonstrated effectiveness in spreadsheet table detection task. We employed the F1 Score as the primary metric to evaluate and compare the performance of different models, as it balances precision and recall, providing a holistic view of model accuracy.
@@ 4.2.1 Dataset
Existing datasets for the Table QA task focus solely on single-table scenarios, leaving a notable gap in performance evaluation for spreadsheets that contain multiple tables. To bridge this gap, we developed a new Spreadsheet QA dataset tailored to the complexities of multi-table environments. We sampled is better 64 spreadsheets from our larger collection and crafted 4-6 questions per spreadsheet, targeting fundamental operations such as searching, comparison, and basic arithmetic. We deliberately excluded questions involving composite operations to maintain clarity and focus in testing specific skills. Each question was paired with an answer, formatted either as a specific cell address or a formula that includes cell addresses, facilitating direct and unambiguous evaluations of the model’s ability to navigate and interpret spreadsheet data. This approach resulted in a comprehensive test dataset comprising 307 items, each a tuple of $(Q,A,S)$, which is detailed in Appendix H.
@@ 4.2.2 Experiment Setup
Given that LLMs have not yet been systematically applied to Spreadsheet QA tasks, we have selected TaPEx and Binder Herzig et al. (2020); Cheng et al. (2022), which are established baselines in the Table QA domain, for comparative evaluation. Since TaPEx and Binder are designed primarily for single-table data, we adapted them for our multi-table context. Initially, our fine-tuned model identifies table regions relevant to each question. These regions are then formatted and fed into the baseline models. In cases where the input exceeds the token limitations of the baseline models, truncation is employed. The accuracy of the answers is assessed based on the correctness of the cell addresses and cell combinations/calculations provided in the answers.
@@ 4.2.3 Experiment Procedure
In this section, we employed the model fine-tuned on the spreadsheet table detection task to conduct QA experiments. The procedure followed the CoS described in section 3.5. Particularly, for instances where the related table was still too large to process effectively, we applied further compression techniques. In cases where tables were exceptionally large and defy effective compression, we utilized a table-splitting algorithm designed to recognize headers and perform strategic concatenation, ensuring that each segment of the split table retains as much contextual integrity as possible. The specifics of this algorithm are detailed in Appendix M.2.
@@ 5.2.2 Ablation Experiment Results
Table 3: Ablation studies on spreadsheet table detection.
@@
Table 3 presents the results of ablation experiments for different modules. The removal of the extraction module led to significantly lower F1 scores, underscoring its critical role in capturing and retaining key structural information. As highlighted in Table 1, this module also achieved the most significant token reduction, confirming its effectiveness. After removing the aggregation module, the F1 score slightly increased. This observation might be attributed to the NFS being more abstract than straightforward numerical representations, which can challenge an LLMs’ ability to interpret them effectively. Despite this, the NFS method offered a significantly high compression rate, enhancing its potential for practical applications and cost control.
@@ 5.3 Spreadsheet QA
Table 4 shows the performance of various models on Spreadsheet QA tasks. We can draw the following conclusions:
1) Effectiveness of the CoS Method: The CoS method we developed significantly boosted the accuracy of models, showing a notable increase of 22% over the baseline GPT4 model. Given the large size of typical spreadsheets, directly inputting entire files often exceeded the token limits of conventional models. The CoS effectively addressed this issue by focusing only on regions relevant to the questions posed, thereby reducing redundant data and enabling more efficient handling of QA tasks on larger spreadsheets.
2) Generalization Capability of the Fine-tuned Model: The model that has been fine-tuned on the spreadsheet table detection task demonstrated robust generalization capabilities across downstream QA tasks. This fine-tuning led to an accuracy improvement of 6%. Moreover, it significantly outperformed the TaPEx and Binder models by 37% and 12%, respectively. This substantial margin highlighted that fine-tuning not only prepared the model to better understand the specific data and structural nuances of spreadsheets but also enhanced its overall comprehension abilities.
Table 4 also shows the influence of our designed split method on QA task performance. It can be seen that using the split algorithm improved accuracy by 3% and 5% on ICL and fine-tuning respectively. Therefore, LLMs can process some tables unable to be processed originally due to the token limitations, which enhances the performance of SpreadsheetLLM.

### v2

@@ Abstract
Spreadsheets are characterized by their extensive two-dimensional grids, flexible layouts, and varied formatting options, which pose significant challenges for large language models (LLMs). In response, we introduce SpreadsheetLLM, pioneering an efficient encoding method designed to unleash and optimize LLMs’ powerful understanding and reasoning capability on spreadsheets. Initially, we propose a vanilla serialization approach that incorporates cell addresses, values, and formats. However, this approach was limited by LLMs’ token constraints, making it impractical for most applications. To tackle this challenge, we develop SheetCompressor, an innovative encoding framework that compresses spreadsheets effectively for LLMs. It comprises three modules: structural-anchor-based compression, inverse index translation, and data-format-aware aggregation. It significantly improves performance in the spreadsheet table detection task, outperforming the vanilla approach by 25.6% in GPT4’s in-context learning setting. Moreover, fine-tuned LLM with SheetCompressor has an average compression ratio of 25×, and achieves a state-of-the-art 78.9% F1 score, surpassing the best existing models by 12.3%. Finally, we propose Chain of Spreadsheet for downstream tasks of spreadsheet understanding and validate it in a new and demanding spreadsheet QA task. We methodically leverage the inherent layout and structure of spreadsheets, demonstrating that SpreadsheetLLM is highly effective across a variety of spreadsheet tasks.
@@
Haoyu Dongthanks:   Corresponding author (hadong@microsoft.com).thanks:   Equal contribution., Jianbo Zhao22footnotemark: 2thanks:   Work during internship at Microsoft., Yuzhang Tian22footnotemark: 233footnotemark: 3, Junyu Xiong22footnotemark: 233footnotemark: 3, Shiyu Xia33footnotemark: 3, Mengyu Zhou, Yun Lin33footnotemark: 3, José Cambronero, Yeye He, Shi Han, Dongmei Zhang Microsoft Corporation
@@ 1 Introduction
We conducted a comprehensive evaluation of our method on a variety of LLMs. Our experiments show that SheetCompressor significantly reduces token usage for spreadsheet encoding by 96%. Moreover, SpreadsheetLLM has shown exceptional performance in spreadsheet table detection, the foundational task of spreadsheet understanding, surpassing the previous SOTA method by 12.3% Dong et al. (2019b). We also applied SpreadsheetLLM to a representative spreadsheet QA task. Inspired by the Chain of Thought (CoT) methodology Zheng et al. (2023); Jiang et al. (2023b), we propose Chain of Spreadsheet (CoS) to decompose spreadsheet reasoning into a table detection-match-reasoning pipeline. It significantly outperformed existing SOTA methods for table QA Liu et al. (2022); Cheng et al. (2022). Our primary contributions are summarized as follows:
@@
  • We fine-tune a variety of cutting-edge LLMs to achieve optimal performance on spreadsheet table detection and demonstrate the high effectiveness of SpreadsheetLLM in accurately understanding complex spreadsheet layouts and structures.
@@
Figure 2: Illustration of the SheetCompressor framework. The original spreadsheet contains two tables, featuring numerous data entries or hierarchical headers. The completed spreadsheet consists of 576 rows and 23 columns, with a vanilla encoding of 61,240 tokens. Initially, we first extract cells using structural anchors, rearranging them into a smaller 24×8 sheet. Subsequently, we perform index-invert, removing empty cells. Finally, we aggregate cells based on data formats, achieving an extremely compact representation of the spreadsheet with only 708 tokens.
@@ 2 Related Work
Spreadsheet representation involves converting the spreadsheets into specific representations for different models. There are various methods for spreadsheet (table) representation. Dong et al. (2019a, b) enhance Mask-RCNN to leverage spatial and visual information in spreadsheets, and Deng et al. (2024) explores the usage of LLMs to evaluate image tables, but it doesn’t work well for spreadsheet images as input to VLMs Xia et al. (2024). To capture sequential semantics in rows and columns, LSTMs are further adopted Nishida et al. (2017); Gol et al. (2019) in row&column directions. Pre-trained LMs Dong et al. (2022) are then proposed to understand spreadsheets Wang et al. (2021). Recent studies Zhang et al. (2023); Li et al. (2023b); Sui et al. (2023) have explored the efficacy of using Markdown, HTML, JSON, and DFLoader for table representation, and these methods are comprehensively summarized in a recent tutorial of Dong and Wang (2024). However, traditional table representations are not well suited to spreadsheets due to their assumption of single table input and absence of explicit cell addresses, as experiments show in Appendix B.
@@
Spreadsheet understanding is enabling for a series of spreadsheet tasks, such as table question answering analysis He et al. (2024); Cheng et al. (2021b, 2022); Jiang et al. (2022); Liu et al. (2022), table extraction Chen and Cafarella (2013, 2014); Li et al. (2024), formula or code generation Chen et al. (2021); Cheng et al. (2021a); Joshi et al. (2024); Chen et al. (2024); Li et al. (2023a), error detection Wang and He (2019); Dou et al. (2016), etc. In this paper, we choose spreadsheet QA, one of the most demanded spreadsheet analysis tasks. It is an extension of the Table QA task in spreadsheet data, with the additional complexity of detecting and matching multiple tables within a spreadsheet.
@@
Related work suggests that the performance of LLMs degrades significantly with long contexts Liu et al. (2024); Xu et al. (2023). Efforts to improve model performance and reduce costs have led to the development of compression techniques for long prompts. Some researchers employ information-theory metrics to filter out redundant information Li (2023); Jiang et al. (2023a). Additionally, specialized models have been proposed to optimize prompt compression Pan et al. (2024). However, these strategies primarily address natural language prompts and may not suit tabular data, potentially leading to considerable structure and data information loss. DBCopilot Wang et al. (2023) enables text-to-SQL conversion on large databases through schema routing. However, due to LLMs’ insufficient ability to understand inherent multi-table layouts and complex table structures that cannot execute queries similar to SQL, schema routing is impractical, restricting the broader application of cutting-edge tabular works Cheng et al. (2022); Li et al. (2023b); Sui et al. (2024) on spreadsheet data.
@@ 3.2 Structural-anchor-based Extraction
where $r_{p}=\left\{Cell_{i,j}\right\}_{i=p,j\in n}$ and $c_{q}=\left\{Cell_{i,j}\right\}_{i\in m,j=q}$. Using these anchor points, our method discards rows and columns that are located more than $k$ cells away from any anchor point, because they rarely serve as table boundaries. The parameter $k$ serves as a threshold to control the scope of neighborhood retention, effectively eliminating areas predominantly filled with homogeneous data that do not contribute to an understanding of the spreadsheet’s layout and structure. We explored the effects of different $k$ values in an ablation study, as detailed in Appendix D.1.
@@ 3.3 Inverted-index Translation
Spreadsheets often contain numerous empty rows, columns, and scattered cells. The standard encoding method, as detailed in Section 3.1, employs a grid-based method that pairs cell addresses with their contents. This approach necessitates recording empty cells to maintain the spreadsheet’s two-dimensional structure, which significantly increases token consumption. Furthermore, cells with identical values are encoded repeatedly, exacerbating token usage.
@@ 3.4 Data-format-aware Aggregation
In spreadsheets, adjacent cells typically share the same data format. As shown in Figure 2 (3), column C records the sell-in billed revenue for different products. Nonetheless, the concrete numerical values are not essential for understanding the structure and semantics of the spreadsheet (although there might be a loss of fine-trained details of exact quantities, e.g., "18,476" and "18,674", this does not impact our comprehension that this column represents revenue). In contrast, the data type is critical for understanding spreadsheets. On one hand, data types represent fundamental semantic properties, such as "time" or "phone number". It motivates us to implement rules to match the value of the cell to different data types. On the other hand, in contrast to detailed numerical values, identical data types may be compressed through clustering, thereby reducing the number of tokens.
@@ 3.5 Chain of Spreadsheet
The query and the identified table section are re-inputted detected tables into the LLM. The model then processes this information to generate an accurate response to the query.
@@ 4.1.1 Dataset
We used the dataset introduced by Dong et al. (2019b), a benchmark dataset of real-world spreadsheets with annotated table boundaries. Due to the complexity and ambiguity of precise address labeling (the Fleiss Kappa on the test set is 0.830), we further implemented the quality improvement pipeline on the test set by five human professionals, as detailed n in Appendix E. To this end, we obtained a highly validated test set containing 188 spreadsheets. Based on the token usage of the vanilla encoding method, we divided the test set into four categories: Small, Medium, Large, and Huge, with a partition of 64:32:70:22. More details are shown in Appendix F. We adopted the Error-of-Boundary 0 (EoB-0) metric for evaluation on 188 spreadsheets with 311 tables. EoB-0 requires exact match of the top, left, bottom, and right boundaries.
@@ 4.1.2 Experiment Setup
To evaluate the performance of SpreadsheetLLM, we chose TableSense-CNN Dong et al. (2019b) as the baseline due to its previously demonstrated effectiveness in the spreadsheet table detection task. We employed the micro F1 Score as the primary metric to evaluate and compare the performance of different models, as it balances precision and recall, providing a holistic view of model accuracy.
@@ 4.2.1 Dataset
Existing datasets for the Table QA task focus solely on single-table scenarios, leaving a notable gap in performance evaluation for spreadsheets that contain multiple tables. To bridge this gap, we developed a new Spreadsheet QA dataset tailored to the complexities of multi-table environments. We sampled 64 spreadsheets from our larger collection and crafted 4-6 questions per spreadsheet, targeting fundamental operations such as searching, comparison, and basic arithmetic. We deliberately excluded questions involving composite operations to maintain clarity and focus in testing specific skills. Each question was paired with an answer, formatted either as a specific cell address or a formula that includes cell addresses, facilitating direct and unambiguous evaluations of the model’s ability to navigate and interpret spreadsheet data. This approach resulted in a comprehensive test dataset comprising 307 items, each a tuple of $(Q,A,S)$, which is detailed in Appendix H.
@@ 4.2.2 Experiment Setup
Given that LLMs have not yet been systematically applied to Spreadsheet QA tasks, we have selected TaPEx and Binder Liu et al. (2022); Cheng et al. (2022), which are established baselines in the Table QA domain, for comparative evaluation. Since TaPEx and Binder are designed primarily for single-table data, we adapted them for our multi-table context. Initially, our fine-tuned model identifies table regions relevant to each question. These regions are then formatted and fed into the baseline models. In cases where the input exceeds the token limitations of the baseline models, truncation is employed. The accuracy of the answers is assessed based on the correctness of the cell addresses and cell combinations/calculations provided in the answers.
@@ 4.2.3 Experiment Procedure
In this section, we employed the model fine-tuned on the spreadsheet table detection task to conduct QA experiments. In the QA experiments, we utilized the fine-tuned model on the table boundary detection task for spreadsheet QA as an ablation to study the generalization capability of the fine- tuned boundary detection model. CoS supports multi-step reasoning during the response generation process. In the QA scenarios, the whole steps include spreadsheet table detection, table structure understanding, table splitting, and sub-table QA. In cases where tables were exceptionally large and defy effective compression, we utilized a table-splitting algorithm designed to recognize headers and perform strategic concatenation, ensuring that each segment of the split table retains as much contextual integrity as possible. The specifics of this algorithm are detailed in Appendix M.2. We categorically classify the model’s responses as either correct or incorrect, and we select accuracy as the evaluation metric.
@@ 5.2.2 Ablation Experiment Results
Table 3: Ablation studies on spreadsheet table detection.
Table 3 presents the results of ablation experiments for different modules. The removal of the extraction module led to significantly lower F1 scores, underscoring its critical role in capturing and retaining key structural information. As highlighted in Table 1, this module also achieved the most significant token reduction, confirming its effectiveness. After removing the aggregation module, the F1 score slightly increased. This observation might be attributed to the NFS being more abstract than straightforward numerical representations, which can challenge an LLM’s ability to interpret them effectively. Despite this, the NFS method offered a significantly high compression rate, enhancing its potential for practical applications and cost control.
@@ 5.3 Spreadsheet QA
Table 4 shows the performance of various models on Spreadsheet QA tasks. We can draw the following conclusions: 1) Effectiveness of the CoS Method: The CoS method with SpreadsheetLLM significantly boosted the accuracy of models, showing a notable increase of 22% over the baseline GPT4 model. Given the large size of typical spreadsheets, directly inputting entire files often exceeded the token limits of conventional models. Some large tables even exceed the token limitation after compression. The CoS effectively addressed this issue by focusing only on regions relevant to the questions posed, improving accuracy by 3% and 5% on ICL and fine-tuning respectively.
In structural-anchor-based extraction, some information loss may occur, which could potentially affect downstream tasks. The CoS method was proposed to address this issue. Based on the overall understanding of the spreadsheet and the table detection, CoS identifies the related table and re-encodes it for downstream tasks without structural-anchor-based extraction. Additionally, the three modules proposed in SpreadsheetLLM are modular and can be freely chosen and used according to different tasks. For instance, the invert-index translation module does not result in any information loss and is almost applicable to all downstream tasks.
In the works of TaPEx and Binder that we referenced, tables are unlike spreadsheets, which can feature a multi-table layout, hundreds of rows (or even more), and formats such as merged cells that complicate the understanding of spatial information. To mitigate these challenges and enhance their capabilities on QA tasks, we have used tables detected in the spreadsheet that were related to the query as inputs. On the other hand, the address information of large tables in spreadsheets is crucial. In SpreadsheetLLM, we have augmented each cell with coordinate information. We speculate that if TaPEx’s pre-training had considered this setting, its performance might have further improved.
2) Generalization Capability of the Fine-tuned Model: The model that has been fine-tuned on the spreadsheet table detection task demonstrated robust generalization capabilities across downstream QA tasks. This fine-tuning on table detection led to an accuracy improvement of 6% on QA. Moreover, it significantly outperformed the TaPEx and Binder models by 37% and 12%, respectively. This substantial margin highlighted that fine-tuning not only prepared the model to better understand the specific data and structural nuances of spreadsheets but also enhanced its overall comprehension abilities.